# Generative Diffusion Models for Resource Allocation in Wireless Networks

### v1

@@ Abstract
This paper proposes a supervised training algorithm for learning stochastic resource allocation policies with generative diffusion models (GDMs). We formulate the allocation problem as the maximization of an ergodic utility function subject to ergodic Quality of Service (QoS) constraints. Given samples from a stochastic expert policy that yields a near-optimal solution to the problem, we train a GDM policy to imitate the expert and generate new samples from the optimal distribution. We achieve near-optimal performance through sequential execution of the generated samples. To enable generalization to a family of network configurations, we parameterize the backward diffusion process with a graph neural network (GNN) architecture. We present numerical results in a case study of power control in multi-user interference networks.
Index Terms:
wireless resource allocation, generative models, diffusion processes, graph neural networks
@@ I Introduction
Most existing formulations and methods for optimal wireless resource allocation, whether classical or learning-based, seek deterministic solutions. In contrast, optimal solutions of many non-convex optimization problems (e.g., power control, scheduling) are inherently probabilistic, as the optimal solution may lie in the convex hull of multiple deterministic policies. By randomizing between multiple deterministic strategies, stochastic policies can achieve better performance by effectively convexifying the problem [1]. This phenomenon is also fundamental in multi-user information theory, where time sharing plays a critical role in achieving optimal performance across various communication channels [2, 3]. In this work, we leverage diffusion models to learn generative solutions to stochastic network resource allocation problems.
@@
A substantial body of the existing literature utilizes GDMs, and GMs in general, for generating domain-specific synthetic data and for data augmentation to enhance the machine-learning models in supervised and reinforcement learning tasks [6, 7]. However, research on the use of GMs for wireless network optimization, and GDMs in particular, is scant [8, 9, 10]. Concurrent works [11, 12, 13, 14] propose generative model solvers for network optimization as a framework to learn solution distributions that concentrate the probability mass around optimal deterministic solutions. The generative (backward) process then converts random noise to high-quality solutions by eliminating the noise introduced in the forward process. However, the problem formulation in the aforementioned studies is deterministic and ignores the probabilistic nature of the optimal solution.
To the best of our knowledge, our work is the first to imitate stochastic expert policies using GDMs. We emphasize the stochastic nature of certain network optimization problems where random solutions are essential for optimality. In our approach, Quality of Service (QoS) near-optimality emerges through the sequential execution of solutions sampled from the optimal generated distribution. Moreover, we use a graph neural network (GNN) architecture as the backbone for the reverse diffusion process to enable learning families of solutions across various network topologies. GNNs not only excel in learning policies from graph-structured data [15, 16] but also exhibit desirable properties such as stability, transferability, and permutation-equivariance.
This paper tackles imitation learning of stochastic wireless resource allocation policies. A GDM policy is trained to match an optimal solution distribution from which an expert policy can sample (see Section II and Section III). We utilize a GNN-parametrization to condition the generative diffusion process directly on the network graphs (see Section IV). We evaluate the proposed GDM policy in a power control setup in multi-user interference networks (see Section V). Our numerical results demonstrate the trained GDM policy closely matches the expert policy over a family of wireless networks.
@@ II Optimal Wireless Resource Allocation
We represent the channel state of a wireless (network) system with a matrix ${\mathbf{H}}\in{\mathcal{H}}\subseteq{\mathbb{R}}^{N\times N}$ and the allocation of corresponding resources with a vector ${\mathbf{x}}\in{\mathcal{X}}\subseteq{\mathbb{R}}^{N}$. Given state ${\mathbf{H}}$, the choice of resource allocation ${\mathbf{x}}$ determines several QoS metrics that we represent with an objective utility $f_{0}:{\mathcal{X}}\times{\mathcal{H}}\mapsto{\mathbb{R}}$ and a constraint utility ${\mathbf{f}}:{\mathcal{X}}\times{\mathcal{H}}\mapsto{\mathbb{R}}^{c}$. We define the optimal resource allocation ${\mathbf{x}}^{*}({\mathbf{H}})$ as an argument that solves the optimization problem,
@@
In (II), we search over stochastic policies $\mathcal{D}_{{\mathbf{x}}}$ that maximize the expected utility $\operatorname{\mathbb{E}}_{\mathcal{D}_{{\mathbf{x}}}}[\,f_{0}({\mathbf{x}}({\mathbf{H}}),{\mathbf{H}})]$ while satisfying the expected constraint $\operatorname{\mathbb{E}}_{\mathcal{D}_{{\mathbf{x}}}}[{\mathbf{f}}\big{(}{\mathbf{x}}({\mathbf{H}}),{\mathbf{H}})]\geq{\mathbf{0}}$ when the resource allocation ${\mathbf{x}}({\mathbf{H}})$ is drawn from the distribution $\mathcal{D}_{{\mathbf{x}}}$. For future reference, we introduce $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{x}}{\,\big{|}\,}{\mathbf{H}})$ to denote a distribution that solves (II). In this distribution, the channel state ${\mathbf{H}}$ is given and resource allocations ${\mathbf{x}}$ are sampled.
The important point here is that the performance of stochastic policies is realizable through time sharing if we allocate resources in a faster time scale than QoS perception. Indeed, if we consider independent resource allocation policies ${\mathbf{x}}_{t}({\mathbf{H}})\sim\mathcal{D}_{{\mathbf{x}}}$ we have that for sufficiently large $T$,
$\frac{1}{T}\sum_{t=1}^{T}f_{0}\big{(}{\mathbf{x}}_{t}({\mathbf{H}}),{\mathbf{H}}\big{)}\approx\operatorname{\mathbb{E}}_{\mathcal{D}_{{\mathbf{x}}}}\Big{[}\,f_{0}\big{(}{\mathbf{x}}({\mathbf{H}}),{\mathbf{H}}\big{)}\,\Big{]},$ (3)
with an analogous statement holding for the constraint utility ${\mathbf{f}}$. Since deterministic policies are particular cases of stochastic policies, we know that ${\text{P}}({\mathbf{H}})\geq\tilde{\text{P}}({\mathbf{H}})$. In practice, it is often the case that ${\text{P}}({\mathbf{H}})\gg\tilde{\text{P}}({\mathbf{H}})$ and for this reason the stochastic formulation in (II) is most often preferred over the deterministic formulation in (II), [1, 2, 3].
@@ II-A Imitation Learning of Stochastic Policies
In this paper, we want to learn to imitate the stochastic policies that solve (II). More to the point, consider a distribution $\mathcal{D}_{{\mathbf{H}}}$ of channel states ${\mathbf{H}}$. For each channel state realization ${\mathbf{H}}$, recall that the solution of (II) is the probability distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{x}}{\,\big{|}\,}{\mathbf{H}})$. Separate from these optimal distributions, we consider a parametric family of conditional distributions $\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}};\boldsymbol{\theta})=\mathcal{D}_{{\mathbf{x}}}({\mathbf{x}}{\,\big{|}\,}{\mathbf{H}};\boldsymbol{\theta})$ in which the channel state ${\mathbf{H}}$ is given, and resource allocation variables are drawn. Our goal is to find the conditional distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}};\boldsymbol{\theta})$ that minimizes the expectation of the conditional KL-divergences $\mathrm{D}_{\mathrm{KL}}\big{(}\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})\;\big{\|}\;\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}};\boldsymbol{\theta})\big{)}$,
@@
In (4), the distributions $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$ are given for all ${\mathbf{H}}$. The conditional distribution $\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}};\boldsymbol{\theta})$ is our optimization variable, which we compare with $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$ through their KL divergence. KL divergences of different channel realizations are averaged over the channel state distribution $\mathcal{D}_{{\mathbf{H}}}$, which is also given. The optimal distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}};\boldsymbol{\theta})=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{x}}{\,\big{|}\,}{\mathbf{H}};\boldsymbol{\theta})$ minimizes the expected KL divergence among distributions that are representable by the parametric family $\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}};\boldsymbol{\theta})$.
@@
A solution of Problem 1 is illustrated in Fig. 1. For a given channel state realization ${\mathbf{H}}$, we show two-dimensional slices of samples of an optimal policy (in blue). As indicated by (2), these samples realize optimal QoS metrics for (II) if executed sequentially (Fig. 2). We train a generative diffusion model (Section III) that generates samples (in orange) that are distributed close to samples of an optimal distribution. When executed sequentially, the learned samples realize QoS metrics close to optimal values (Fig. 2). Neither the optimal distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})\mathcal{D}_{{\mathbf{H}}}=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{x}}{\,\big{|}\,}{\mathbf{H}})\mathcal{D}_{{\mathbf{H}}}$ nor the parametric distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}};\boldsymbol{\theta})=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{x}}{\,\big{|}\,}{\mathbf{H}};\boldsymbol{\theta})$ is computed.
Remark 1
Most learning approaches to allocating resources in wireless systems contend with the deterministic policy formulation in (II), e.g., [15, 17, 22, 10, 11, 12, 14, 13]. This is due in part to the use of deterministic learning parameterizations [15, 17, 22] but even recent contributions that propose diffusion models, do so for deterministic policies [10, 11, 12, 14, 13]. This is a well-known limitation that has motivated, e.g., state-augmented algorithms that leverage dual gradient descent dynamics to randomize policy samples [16]. Our paper is the first to develop and demonstrate imitation of stochastic policies with generative diffusion models.
@@ III Policy Generative Models
GDMs involve a forward and a backward diffusion process. The forward process defines a Markov chain of diffusion steps to progressively add random noise to data. For a given ${\mathbf{H}}$ and a data sample ${\mathbf{x}}_{0}={\mathbf{x}}_{0}({\mathbf{H}})$ drawn from the expert distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$, the forward chain follows
$q({\mathbf{x}}_{t}\,|\,{\mathbf{x}}_{0};{\mathbf{H}})=\mathcal{N}({\mathbf{x}}_{t};\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0},(1-\bar{\alpha}_{t}){\mathbf{I}}),$ (5)
where $\bar{\alpha}_{t}:=\prod_{i=1}^{t}\alpha_{i}$, $\alpha_{t}:=1-\beta_{t}$, and $\beta_{t}$ is a monotonically increasing noise schedule, e.g., linear. For a sufficiently large $T$, (5) converts the data sample ${\mathbf{x}}_{0}$ into a sample that is approximately isotropic Gaussian distributed, i.e., ${\mathbf{x}}_{T}\approx\mathcal{N}({\mathbf{0}},{\mathbf{I}})$.
The reverse process of (5) can be approximated as a chain of Gaussian transitions with some fixed variance $\sigma^{2}_{t}{\mathbf{I}}$ and a parametrized mean $\boldsymbol{\mu}_{\boldsymbol{\theta}}$,
$p_{\boldsymbol{\theta}}({\mathbf{x}}_{t-1}\,|\,{\mathbf{x}}_{t};{\mathbf{H}})=\mathcal{N}\big{(}{\mathbf{x}}_{t-1};\boldsymbol{\mu}_{\boldsymbol{\theta}}({\mathbf{x}}_{t},t;{\mathbf{H}}),\sigma^{2}_{t}{\mathbf{I}}\big{)}.$ (6)
A backward diffusion process samples ${\mathbf{x}}_{T}\sim\mathcal{N}({\mathbf{0}},{\mathbf{I}})$ and iteratively runs the backward chain in (6) for $t=T,\ldots,1$. Reparametrizing (5) as ${\mathbf{x}}_{t}({\mathbf{x}}_{0},\boldsymbol{\epsilon})=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ [4], we equate sampling ${\mathbf{x}}_{t-1}\sim p_{\boldsymbol{\theta}}(.\,|\,{\mathbf{x}}_{t};{\mathbf{H}})$ to computing
${\mathbf{x}}_{t-1}=\frac{1}{\sqrt{\alpha}_{t}}\left({\mathbf{x}}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\boldsymbol{\theta}}({\mathbf{x}}_{t},t;{\mathbf{H}})\right)+\sigma_{t}{\mathbf{w}},$ (7)
where ${\mathbf{w}}\sim\mathcal{N}({\mathbf{0}},{\mathbf{I}})$, and $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}({\mathbf{x}}_{t},t;{\mathbf{H}})$ predicts the noise $\boldsymbol{\epsilon}$ added to ${\mathbf{x}}_{0}\sim\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$ from noisy sample ${\mathbf{x}}_{t}$ at timestep $t$.
@@
${\mathcal{L}}_{\mathrm{GDM}}(\boldsymbol{\theta})=\operatorname{\mathbb{E}}_{{\mathbf{x}}_{0},t,{\mathbf{H}},\boldsymbol{\epsilon}}\,\omega(t)\!\left\|\boldsymbol{\epsilon}_{\boldsymbol{\theta}}\big{(}{\mathbf{x}}_{t}({\mathbf{x}}_{0},\boldsymbol{\epsilon}),t;{\mathbf{H}}\big{)}-\boldsymbol{\epsilon}\right\|^{2}\!.$ (8)
In (8), $\omega(t)$ is a time-dependent weighting function and the expectation is over random timesteps $t\sim\mathrm{Uniform}([1,T])$, Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}\left({\mathbf{0}},{\mathbf{I}}\right)$, expert (data) samples ${\mathbf{x}}_{0}\sim\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$, and conditioning networks ${\mathbf{H}}\sim\mathcal{D}_{{\mathbf{H}}}$.
The DDPM loss in (8) is a variational upper bound on the expected KL divergence loss in (4), which becomes tight when $\boldsymbol{\theta}=\boldsymbol{\theta}^{*}$. Thus, running (7) with optimal parametrization $\boldsymbol{\epsilon}_{\boldsymbol{\theta}^{*}}$ for a given ${\mathbf{H}}$ generates samples from the expert conditional distribution, i.e., ${\mathbf{x}}_{0}\sim\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}};\boldsymbol{\theta}^{*})=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}};\boldsymbol{\theta})\approx\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$.

### v3

@@ Abstract
This paper proposes a supervised training algorithm for learning stochastic resource allocation policies with generative diffusion models (GDMs). We formulate the allocation problem as the maximization of an ergodic utility function subject to ergodic Quality of Service (QoS) constraints. Given samples from a stochastic expert policy that yields a near-optimal solution to the constrained optimization problem, we train a GDM policy to imitate the expert and generate new samples from the optimal distribution. We achieve near-optimal performance through the sequential execution of the generated samples. To enable generalization to a family of network configurations, we parameterize the backward diffusion process with a graph neural network (GNN) architecture. We present numerical results in a case study of power control.
@@ I Introduction
Most existing formulations and methods for optimal wireless resource allocation, whether classical or learning-based, seek deterministic solutions. In contrast, optimal solutions of many non-convex optimization problems (e.g., power control, scheduling) are inherently probabilistic, as the optimal solution may lie in the convex hull of multiple deterministic policies. By randomizing between multiple deterministic strategies, stochastic policies can achieve better performance by effectively convexifying the problem [1]. This phenomenon is also fundamental in multi-user information theory, where time sharing plays a critical role in achieving optimal performance across various communication channels [2, 3, 4]. In this work, we leverage diffusion models to learn generative solutions to stochastic network resource allocation problems.
@@
A substantial body of the existing literature utilizes GDMs, and GMs in general, for generating domain-specific synthetic data and for data augmentation to enhance the machine-learning models in supervised and reinforcement learning tasks [7, 8]. Yet, research on the use of GMs for wireless network optimization, and GDMs in particular, is scant [9, 10, 11, 12, 13]. Concurrent works [14, 15, 16, 17, 18, 19] propose generative model solvers for network optimization as a framework to learn solution distributions that concentrate the probability mass around optimal deterministic solutions. The generative process then converts random noise to high-quality solutions by eliminating the noise introduced in the forward process. However, the problem formulation in the aforementioned studies is deterministic and ignores the probabilistic nature of the optimal solution.
Our work is one of the first to imitate stochastic expert policies using GDMs. We emphasize the stochastic nature of certain network optimization problems where random solutions are not only essential for optimality but also are realized by leveraging iterative dual domain algorithms. In our approach, Quality of Service (QoS) near-optimality emerges through the sequential execution of solutions sampled from the optimal generated distribution. Moreover, we use a graph neural network (GNN) architecture as the backbone for the reverse diffusion process to enable learning families of solutions across network topologies. GNNs not only excel in learning policies from graph-structured data [20, 21, 22, 23] but also exhibit desirable properties such as stability, transferability and permutation-equivariance [24, 25].
This paper tackles imitation learning of stochastic wireless resource allocation policies. A GDM policy is trained to match an optimal solution distribution to a constrained optimization problem from which an expert policy can sample (Section II and Section III). We utilize a GNN-parametrization to condition the generative diffusion process directly on the network graphs (Section IV). We evaluate the proposed GDM policy in a power control setup and demonstrate that the trained GDM policy closely matches the expert policy over a family of wireless networks (Section V).
@@ II Optimal Wireless Resource Allocation
We represent the channel state of a wireless (network) system with a matrix ${\mathbf{H}}\in{\mathcal{H}}\subseteq{\mathbb{R}}^{N\times N}$ and the allocation of corresponding resources with a vector ${\mathbf{x}}\in{\mathcal{X}}\subseteq{\mathbb{R}}^{N}$. Given $\mathbf{H}$}, the choice of resource allocation $\mathbf{x}$} determines several QoS metrics that we represent with an objective utility $f_{0}:{\mathcal{X}}\times{\mathcal{H}}\mapsto{\mathbb{R}}$ and a constraint utility ${\mathbf{f}}:{\mathcal{X}}\times{\mathcal{H}}\mapsto{\mathbb{R}}^{c}$. We define an optimal resource allocation ${\mathbf{x}}^{*}({\mathbf{H}})$ as the argument that solves the constrained optimization problem,
@@
In (II), we search over stochastic policies $\mathcal{D}_{{\mathbf{x}}}$ that maximize the expected utility $\operatorname{\mathbb{E}}_{\mathcal{D}_{{\mathbf{x}}}}[\,f_{0}({\mathbf{x}}({\mathbf{H}}),{\mathbf{H}})]$ while satisfying the expected constraint $\operatorname{\mathbb{E}}_{\mathcal{D}_{{\mathbf{x}}}}[{\mathbf{f}}\big({\mathbf{x}}({\mathbf{H}}),{\mathbf{H}})]\geq{\mathbf{0}}$ when the resource allocation ${\mathbf{x}}({\mathbf{H}})$ is drawn from the distribution $\mathcal{D}_{{\mathbf{x}}}$. For future reference, we introduce $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{x}}{\,\big|\,}{\mathbf{H}})$ to denote a distribution that solves (II). In $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$, the channel state $\mathbf{H}$} is given and allocations $\mathbf{x}$} are sampled.
The important point here is that the performance of stochastic policies is realizable through time sharing if we allocate resources in a faster time scale than QoS perception. Indeed, if we consider independent resource allocation policies ${\mathbf{x}}_{\tau}({\mathbf{H}})\sim\mathcal{D}_{{\mathbf{x}}}$ we have that for sufficiently large $T$,
$\frac{1}{T}\sum_{\tau=1}^{T}f_{0}\big({\mathbf{x}}_{\tau}({\mathbf{H}}),{\mathbf{H}}\big)\approx\operatorname{\mathbb{E}}_{\mathcal{D}_{{\mathbf{x}}}}\Big[\,f_{0}\big({\mathbf{x}}({\mathbf{H}}),{\mathbf{H}}\big)\,\Big],$ (3)
with an analogous statement holding for the constraint utility $\mathbf{f}$}. Since deterministic policies are particular cases of stochastic policies, we know that ${\text{P}}({\mathbf{H}})\geq\tilde{\text{P}}({\mathbf{H}})$. In practice, it is often the case that ${\text{P}}({\mathbf{H}})\gg\tilde{\text{P}}({\mathbf{H}})$ and for this reason, the stochastic formulation in (II) is most often preferred over the deterministic formulation in (II), [1, 2, 3].
@@ II-A Imitation Learning of Stochastic Policies
In this paper, we want to learn to imitate the stochastic policies that solve (II). Consider a distribution $\mathcal{D}_{{\mathbf{H}}}$ of channel states $\mathbf{H}$}. For each realization $\mathbf{H}$}, recall that the solution of (II) is the probability distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{x}}{\,\big|\,}{\mathbf{H}})$. Separate from these optimal distributions, we consider a parametric family of conditional distributions $\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}};\boldsymbol{\theta})=\mathcal{D}_{{\mathbf{x}}}({\mathbf{x}}{\,\big|\,}{\mathbf{H}};\boldsymbol{\theta})$ in which the channel state $\mathbf{H}$} is given, and resource allocation variables are drawn. Our goal is to find the conditional distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}};\boldsymbol{\theta})$ that minimizes the expectation of the conditional KL-divergences $\mathrm{D}_{\mathrm{KL}}\big(\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})\;\big\|\;\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}};\boldsymbol{\theta})\big)$,
@@
In (4), the distributions $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$ are given for all $\mathbf{H}$}. The conditional distribution $\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}};\boldsymbol{\theta})$ is our optimization variable, which we compare with $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$ through their KL divergence. KL divergences of different channel realizations are averaged over the channel state distribution $\mathcal{D}_{{\mathbf{H}}}$, which is also given. The optimal distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}};\boldsymbol{\theta})=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{x}}{\,\big|\,}{\mathbf{H}};\boldsymbol{\theta})$ minimizes the expected KL divergence among those representable by the parametric family $\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}};\boldsymbol{\theta})$.
@@
A solution of Problem 1 is illustrated in Fig. 2. For a given channel state realization $\mathbf{H}$}, we show two-dimensional slices of samples of an optimal policy (in blue). As indicated by (1), these samples realize optimal QoS metrics for (II) if executed sequentially (Fig. 1). We train a generative diffusion model (Section III) that generates samples (in orange) that are distributed close to samples of an optimal distribution. When executed sequentially, the learned samples realize QoS metrics close to optimal values (Fig. 1). We underscore that neither the optimal distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})\mathcal{D}_{{\mathbf{H}}}=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{x}}{\,\big|\,}{\mathbf{H}})\mathcal{D}_{{\mathbf{H}}}$ nor the parametric distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}};\boldsymbol{\theta})=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{x}}{\,\big|\,}{\mathbf{H}};\boldsymbol{\theta})$ is computed.
II-B Learning in the Dual Domain & Policy Randomization
Most learning approaches to allocating resources in wireless systems contend with the deterministic policy formulation in (II), e.g., [21, 26, 31, 11, 14, 15, 17, 16]. This is due in part to the use of deterministic learning parameterizations [21, 26, 31] but even recent contributions that propose diffusion models do so for deterministic policies [11, 14, 15, 17, 16, 18]. This is a well-known limitation that has motivated, e.g., state-augmented algorithms that leverage dual gradient descent dynamics to randomize policy samples [22, 23]. These algorithms generate trajectories of primal and dual iterates by operating on a convex hull relaxation of the Lagrangian for the original problem and iteratively solving a sequence of Lagrangian maximization subproblems. Each subproblem is an unconstrained, deterministic problem to which regular learning methods apply, and near-optimality and feasibility guarantees are established neither for individual primal iterates nor their averages, but only for the sequential execution of the generated policy iterates.
A shortcoming of state-augmented algorithms—and iterative dual domain algorithms in general—is that they incur a transient period where suboptimal policies are executed. Reducing the length of this transient period typically requires larger step sizes, which in turn introduces a trade-off with respect to solution optimality. Learning a generative model to sample from the stationary (optimal) policy distribution emerges as a promising approach for overcoming this trade-off. To the best of our knowledge, our paper is the first to develop and demonstrate imitation of stochastic policies that solve a constrained optimization problem with generative diffusion models.
@@ III Policy Generative Models
GDMs involve a forward and a backward diffusion process. The forward process defines a Markov chain of diffusion steps to progressively add random noise to data. For a given $\mathbf{H}$} and a data sample ${\mathbf{x}}_{0}={\mathbf{x}}({\mathbf{H}})$ drawn from the expert distribution $\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$, the forward chain follows
$q({\mathbf{x}}_{k}\,|\,{\mathbf{x}}_{0};{\mathbf{H}})=\mathcal{N}({\mathbf{x}}_{k};\sqrt{\bar{\alpha}_{k}}{\mathbf{x}}_{0},(1-\bar{\alpha}_{k}){\mathbf{I}}),$ (5)
where $\bar{\alpha}_{k}:=\prod_{i=1}^{k}\alpha_{i}$, $\alpha_{k}:=1-\beta_{k}$, and $\beta_{k}$ is a monotonically increasing noise schedule, e.g., linear. For a sufficiently large $K$, (5) converts the data sample ${\mathbf{x}}_{0}$ into a sample that is approximately isotropic Gaussian distributed, i.e., ${\mathbf{x}}_{K}\approx\mathcal{N}({\mathbf{0}},{\mathbf{I}})$.
The reverse process of (5) is approximated by a chain of Gaussian transitions with a parametrized mean $\boldsymbol{\mu}_{\boldsymbol{\theta}}$ and fixed variance $\sigma^{2}_{k}{\mathbf{I}}$,
$p_{\boldsymbol{\theta}}({\mathbf{x}}_{k-1}\,|\,{\mathbf{x}}_{k};{\mathbf{H}})=\mathcal{N}\big({\mathbf{x}}_{k-1};\boldsymbol{\mu}_{\boldsymbol{\theta}}({\mathbf{x}}_{k},k;{\mathbf{H}}),\sigma^{2}_{k}{\mathbf{I}}\big).$ (6)
A backward diffusion process samples ${\mathbf{x}}_{K}\sim\mathcal{N}({\mathbf{0}},{\mathbf{I}})$ and iteratively runs the backward chain in (6) for $k=K,\ldots,1$. With reparametrization of (5) as ${\mathbf{x}}_{k}({\mathbf{x}}_{0},\boldsymbol{\epsilon})=\sqrt{\bar{\alpha}_{k}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{k}}\boldsymbol{\epsilon}$ [5], sampling ${\mathbf{x}}_{k-1}\sim p_{\boldsymbol{\theta}}(.\,|\,{\mathbf{x}}_{k};{\mathbf{H}})$ amounts to updating
${\mathbf{x}}_{k-1}=\frac{1}{\sqrt{\alpha}_{k}}\left({\mathbf{x}}_{k}-\frac{\beta_{k}}{\sqrt{1-\bar{\alpha}_{k}}}\boldsymbol{\epsilon}_{\boldsymbol{\theta}}({\mathbf{x}}_{k},k;{\mathbf{H}})\right)+\sigma_{k}{\mathbf{w}},$ (7)
where ${\mathbf{w}}\sim\mathcal{N}({\mathbf{0}},{\mathbf{I}})$, and $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}({\mathbf{x}}_{k},k;{\mathbf{H}})$ predicts the noise $\boldsymbol{\epsilon}$ added to ${\mathbf{x}}_{0}\sim\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$ from noisy sample ${\mathbf{x}}_{k}$ at timestep $k$.
@@
${\mathcal{L}}_{\mathrm{GDM}}(\boldsymbol{\theta})=\operatorname{\mathbb{E}}_{{\mathbf{x}}_{0},k,{\mathbf{H}},\boldsymbol{\epsilon}}\,\omega(k)\!\left\|\boldsymbol{\epsilon}_{\boldsymbol{\theta}}\big({\mathbf{x}}_{k}({\mathbf{x}}_{0},\boldsymbol{\epsilon}),k;{\mathbf{H}}\big)-\boldsymbol{\epsilon}\right\|^{2}\!.$ (8)
In (8), $\omega(k)$ is a time-dependent weighting function and the expectation is over random timesteps $k\sim\mathrm{Uniform}([1,K])$, Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}\left({\mathbf{0}},{\mathbf{I}}\right)$, expert (data) samples ${\mathbf{x}}_{0}\sim\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$, and conditioning networks ${\mathbf{H}}\sim\mathcal{D}_{{\mathbf{H}}}$.
We note that the DDPM loss in (8) is a variational upper bound on the expected KL divergence loss in (4), which becomes tight when $\boldsymbol{\theta}=\boldsymbol{\theta}^{*}$. Thus, running (7) with optimal parametrization $\boldsymbol{\epsilon}_{\boldsymbol{\theta}^{*}}$ for a given $\mathbf{H}$} generates samples from the expert conditional distribution, i.e., ${\mathbf{x}}_{0}\sim\mathcal{D}_{{\mathbf{x}}}({\mathbf{H}};\boldsymbol{\theta}^{*})=\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}};\boldsymbol{\theta})\approx\mathcal{D}_{{\mathbf{x}}}^{*}({\mathbf{H}})$.